# 100,000 Crab giant pulses at 215 MHz detected with an SKA-Low prototype station

### v1

@@ 4.2 Flux Density Calibration
where and are observing bandwidth and integration time respectively. is the System Equivalent Flux Density (SEFD) for Stokes I polarisation (hence and this factor is not included under the square root). For each dataset, was calculated using the SKA-Low-sensitivity calculator (2022PASA...39...15S)666https://github.com/marcinsokolowski/station_beam, which provides very accurate SEFD values as verified by extensive tests during the EDA2 and AAVS2 commissioning and verification stages (2022JATIS...8a1010W; 2021ecap.confE...1S). The mean value during the Crab pulsar transit at elevation 41°was Jy (for reference EDA2 SEFD at zenith is 2500 Jy). Hence, the expected noise was Jy for the observing parameters Hz (=31.25 MHz), sec (1 ms).
@@ 5.1 Fluence distribution
Approximately 85% of the detected GPs occur at phase of the MP (the remaining 15% at phase with IP), which is in very good agreement with 86% measured at similar frequencies with MWA and 84% at 732 MHz (2017ApJ...851...20M). It is also very close to 87% at 1200 MHz (2012ApJ...760...64M).
@@ 5.2 Scatter broadening and DM variations
Scattering time and DM as a function of time are shown in Figure 3, where DM measured by Jodrell Bank (1993MNRAS.265.1003L)888http://www.jb.man.ac.uk/~pulsar/crab/crab2.txt is also plotted as a reference. A systematic constant offset between DM measured in this work and Jodrell Bank measurements is noticeable (see also Figure 16 and discussion in B). Especially, at MJDs between the start and 60680 days, before the increase in scattering time begun, which could affect our low-frequency timing-based DM measurements. Such DM offsets are not uncommon for measurements performed at different frequencies (this work at 215 MHz and Jodrell Bank in L-band), and may result from chromatic DM (see, for example2019ApJ...882..133K; 2016ApJ...817...16C). However, this does not affect our analysis.
@@
The DM was measured using standard pulsar timing performed with the TEMPO2 software package using frequency dependent pulse profiles. These measurements may be impacted (overestimate DM) by the changes of the average pulsar profile due to scattering - mainly by shifting the peak and changing the shape of the average pulse profile. Therefore, as a cross-check the DM was also measured by fitting profiles of individual GPs and finding DM maximising their peak flux density (i.e. SNR). The comparison of the DM measured with this method and timing is shown in Fig. 16 in B, where we also discuss other tested methods of measuring DM from individual GPs. The DM values measured from individual GPs are similar to the results of the timing analysis, and confirm that to the first order latter are correct and the increase in DM is of physical origin (not an artefact of analysis). Although, during high scatter broadening period (between MJD 60700 and 60740), the increase in DM may be slightly lower than shown by the red filled circles in Figure 3, the DM values lower bound by the well-established Jodrell Bank measurements (red stars). This ensures that our further analysis is valid to within a factor of 2 (in the worst case).
@@
There is a strong positive correlation between DM and scattering time (Fig. 5), and as the scattering time gets higher DM also increases and vice versa (the Pearson correlation coefficient 0.7 and the p-value ). However, we note that this correlation may be slightly exaggerated by the earlier-mentioned impact of scattering on timing-based DM measurements. A similarly strong correlation of scatter broadening time and DM (correlation coefficient 0.560.01) based on normal pulses was observed in Crab pulsar at 610 MHz by 2018MNRAS.479.4216M, and at 350 MHz by 2019MNRAS.483.1224D. Hence, our measurements are consistent with the earlier observations of scatter broadening and DM correlations at higher frequencies. Nevertheless, we note that the data are very limited and our low-frequency observations are unique (especially in the context of new generation of instruments).
@@
The lag correlation of DM and scatter broadening time variations yielded a lag consistent with zero (to within a daily cadence of our observations). This is consistent with the results reported by 2018MNRAS.479.4216M, but significantly differs from the 30 days lag observed by 2008A&A...483...13K. It seems that the DM/ increase event observed by 2008A&A...483...13K was caused by a plasma screen with a different structure and/or geometry of the Earth, pulsar, screen.
@@
Finally, we have also compared scatter broadening times with earlier observations. Thus, the measured scatter broadening times were scaled to 610 MHz with a scaling factor as measured for the Crab pulsar (2019ApJ...874..179K; 2007ApJ...665..618B; 2017ApJ...851...20M). The resulting distribution is very similar and only slightly shifted towards lower scatter-broadening times with respect to the distribution in Figure 3 in 2018MNRAS.479.4216M. This indicates that the scattering conditions were likely similar during both events, and possibly caused by similar plasma structures.
@@
One of the most interesting aspects originally noticed during the first observations in December 2024, were large variations in the numbers of detected GP between 1-hour observations on different days (i.e. the rate of detectable GPs varied). During the first night of observations (2024-12-14) about 1400 GPs were detected by PRESTO, while on 2024-12-25 it reached 3000 to drop down to around 250 in the middle of February 2025 (Fig. 4). Interestingly, it started to increase again at the end of February 2025 to reach 2000 at the end of March (on 2025-03-27). The number of detected GPs was strongly correlated with the scatter broadening time (compare Figures 3 and 4, and see correlation in Figure 6).
The number of detected GPs as a function of scattering time has been modelled (black curve in Figure 6) using the scatter broadening time dependent fluence distribution (the fitted linear function in Figure 7), and functional form of pulse profile. Equation 15 in A was used to model giant pulse profiles as a function of time for different values of scatter broadening time with other parameters fixed (, , and ms). As discussed in Section 5.3 the fluence distribution of GPs depends on scatter broadening time. Hence, we used the linear dependence as fitted in Figure 7 to calculate normalisation of the fluence distribution for a specific . Then the expected number of GPs for particular value of was calculated as an integral of the fluence distribution above the minimal fluence . This minimal fluence was calculated as the fluence corresponding to the SNR threshold () used by the detection software PRESTO (SNR=5, i.e. peak flux density = 5230 Jy ms). The number of GPs as a function of and the results of our modelling are shown as a black curve in Figure 6. We note that this is only an approximation, because as discussed earlier by using convolution with boxcar function, PRESTO effectively implements a fluence detector, and this has not been reflected in our model. However, even this simple model reproduces the observed dependence very well. This gives us confidence that the variations in number of detected GPs can be explained by the up and down variations of the fluence distribution caused by refractive scintillation, which is also supported by the weeks timescale of these variations. Similar, variations in GP rates were observed by 1995ApJ...453..433L in their largest ever recorded sample of GPs, which they also attributed to propagation effects. The explanation in terms of refractive scintillation is discussed in Section 6.1.2.
@@ 5.4 Flux density, SNR and spectral density distributions
Using the calibrated peak flux density and the distance to the pulsar D=2 kpc (2023ApJ...952..161L; 2008ApJ...677.1201K), the peak spectral luminosity was calculated according to the following equation:
where is the peak flux density of the pulse. The resulting distribution is shown in the right panel of Figure 8. The maximum values are Lerg/s/Hz, which is consistent with the Crab spectral luminosities reported in literature (e.g. compare with Fig. 3 in 2022NatAs...6..393N or Fig. 1 in 2018NatAs...2..865K). These maximum spectral luminosities are a 2 – 3 orders of magnitude lower than the weakest repeating FRBs. However, assuming that there is enough energy in the system, the wait time for erg/s/Hz pulse is relatively short (of the order of 50,000 years). Hence, extra-galactic Crab-like pulsars may be responsible for weak repeating FRBs from the local Universe (see discussion in Section 6.2.1). However, the observed spectral luminosities are more than 10 orders of magnitude lower than typical or brightest FRBs, making it highly unlikely that such FRBs can be produced by objects similar to Crab pulsar (see discussion in Section 6.2). Hence, there is simply not enough energy available. Additionally, based on our spectral luminosity distribution, the waiting time for such bright pulses (erg/s/Hz) is more than 10 orders of magnitude longer than the Hubble time.
@@ 6.1 Modelling DM and variations
We consider two distances between Earth and the screen (): (i) standard case midway (1000 pc) between the Earth and pulsar, and (ii) close to the pulsar (some 1900 pc from Earth) and consistent with the Crab Nebula or more broadly Perseus arm of the galaxy (explanation will be provided in Section 6.1.2). These two cases correspond to the ratio x= (Fig. 10) equal to 1 for (i) and 20 for (ii) respectively. For x=1 the velocity of the scintillation pattern with respect to observer on Earth is approximately equal to the velocity of the pulsar (eq. 10), while it can be higher for due to “lever-arm” argument (see equation 10 or discussion in 1995ApJ...451..717G).
@@ 6.1.1 Modelling DM variations
The characteristic spatial scale of the plasma blob resulting from the fit is pc, and pc is consistent with the cloud size R pc estimated above using the mean values. This value leads to peak electron density 2600 e-cm-3. As an additional check, we performed a standard DM structure function analysis (for description see Section 5.2 in 2018MNRAS.479.4216M) and obtained a characteristic timescale of 12.9 days, which corresponds to pc (when multiplied by 120 km/s), which agrees nearly perfectly with the value.
Hence, all tested methods lead to consistent results, and the resulting parameters are physically plausible. Despite the different epochs of observations, frequencies etc., they are also in-line with the studies by 2018MNRAS.479.4216M; 2008A&A...483...13K; 2011MNRAS.410..499G as summarised in Table 1. This suggests that similar scintillation/scattering events due to plasma screens of the size pc and electron densities of the order of – are relatively common for the Crab pulsar. The observed electron densities are consistent with the filamentary structures inside the Crab Nebula (e.g. 1957PASP...69..227O; 1982ApJ...258....1F).
@@
The baseline level of scattering time in this work was around 0.1 ms which corresponds to 2.5 ms at 215 MHz. However, sudden drops down to 0 ms were also observed (see Figure 2 in 2018MNRAS.479.4216M). Since we are not sure if these were valid measurements of extremely low scattering time we used the baseline level as a reference.
@@ 6.1.2 Effects of refractive scintillation
The long observed timescale 70 days of the variations in GP rate and fluence distribution is a manifestation of intensity modulations caused by refractive scintillation. Moreover, estimates of the size pc of the plasma cloud agree very well with the large-scale structures required by refractive scintillation (see, for example 1999ApJ...514..272B). The refractive scintillation effects on Crab GPs and average profiles have been observed and discussed before (e.g. 1995ApJ...453..433L, 1990MNRAS.244...68R, and 2011MNRAS.410..499G). Interestingly, in a recent paper, 2024ApJ...973...87D discuss the possibility of using GP rate as an ISM probe, which is exactly what transpired from our observations.
Under this hypothesis and using our equation 9 and equation 4.45 from 2012hpa..book.....L (see also equation 4.1 in 1986MNRAS.220...19R), we obtain:
@@
Additionally, the scatter broadening time is related to by the equation (e.g. (2012puas.book.....L)) :
@@
Typically, scintillation analysis (e.g. 2014ApJ...791L..32B), uses auto-correlation function (ACF) of dynamic spectrum containing scintles to calculate scintillation time and de-correlation bandwidth as respectively the full-width at 1/e and half-maximum half-width of the 2D Gaussian fitted to the correlation peak in ACF. In this analysis the fluence normalisation or number of GPs were sampled daily in 1 hour observations. Hence, long-timescale dynamic spectrum is not available, and we can only use 1D temporal dependence of the number of GPs or normalisation of the fluence distribution (). Hence, in order to be consistent with this convention, we estimate by fitting Gaussian to the peak of the normalisation of the fluence distribution (not shown here but very similar to Figure 4) as a function of time, and calculate 51 days as full width at 1/e of the peak.
In order to calculate using equation 8, the distance to the screen and are also required. The additional constraints are imposed by equation 9 (above) and the following equation (equation 3 in 2014ApJ...791L..32B):
@@
The first solution 0.05 pc could potentially be related to the inner edge of the solar system (distance 0.01 pc), but this does not have convincing physical explanation. On the other hand, the second solution (1835 pc) places the screen close to the pulsar, and the distance matches the distance to the Perseus spiral arm of the Milky Way galaxy, where the Crab pulsar is located. Hence, the scattering screen could be placed somewhere at the edge of the Perseus arm as seen along the LoS from Earth, which was also proposed by earlier studies (e.g. 1971ApJ...166..513C). Alternatively, given the accuracy of this simplistic approach, and uncertainties of the distance to the pulsar, velocities of the screen and pulsar it is possible that the scattering material corresponds to the expanding Crab Nebula surrounding the pulsar. We note that solar elongations of Crab pulsar during the observing period were between 180°and 74°- sufficiently large to make any contributions from the solar wind negligible.

### v2

@@ 4.2 Flux Density Calibration
where and are observing bandwidth and integration time respectively. is the System Equivalent Flux Density (SEFD) for Stokes I polarisation (hence and this factor is not included under the square root). For each dataset, was calculated using the SKA-Low-sensitivity calculator 2022PASA…39…15S (79)666https://github.com/marcinsokolowski/station_beam, which provides very accurate SEFD values as verified by extensive tests during the EDA2 and AAVS2 commissioning and verification stages 2022JATIS…8a1010W (89, 82). The mean value during the Crab pulsar transit at elevation 41°was Jy (for reference EDA2 SEFD at zenith is 2500 Jy). Hence, the expected noise was Jy for the observing parameters Hz (=31.25 MHz), sec (1 ms).
@@ 5.1 Fluence distribution
Approximately 85% of the detected GPs occur at phase of the MP (the remaining 15% at phase with IP), which is in very good agreement with 86% measured at similar frequencies with MWA and 84% at 732 MHz 2017ApJ…851…20M (63). It is also very close to 87% at 1200 MHz 2012ApJ…760…64M (64).
@@ 5.2 Scatter broadening and DM variations
Scattering time and DM as a function of time are shown in Figure 3, where DM measured by Jodrell Bank 1993MNRAS.265.1003L (55)888http://www.jb.man.ac.uk/~pulsar/crab/crab2.txt is also plotted as a reference. A systematic constant offset between DM measured in this work and Jodrell Bank measurements is noticeable (see also Figure 17 and discussion in B). Especially, at MJDs between the start and 60680 days, before the increase in scattering time begun, which could affect our low-frequency timing-based DM measurements. Such DM offsets are not uncommon for measurements performed at different frequencies (this work at 215 MHz and Jodrell Bank in L-band), and may result from chromatic DM (see, for example2019ApJ…882..133K (42, 16)). However, this does not affect our analysis.
@@
The DM was measured using standard pulsar timing performed with the TEMPO2 software package using frequency dependent pulse profiles. These measurements may be impacted (overestimate DM) by the changes of the average pulsar profile due to scattering - mainly by shifting the peak and changing the shape of the average pulse profile. Therefore, as a cross-check the DM was also measured by fitting profiles of individual GPs and finding DM maximising their peak flux density (i.e. SNR). The comparison of the DM measured with this method and timing is shown in Fig. 17 in B, where we also discuss other tested methods of measuring DM from individual GPs. The DM values measured from individual GPs are similar to the results of the timing analysis, and confirm that to the first order the latter are correct and the increase in DM is of physical origin (not an artefact of analysis). Although, during high scatter broadening period (between MJD 60700 and 60740), the increase in DM may be slightly lower than shown by the red filled circles in Figure 3, the DM values are lower bound by the well-established Jodrell Bank measurements (red stars). This ensures that our further analysis is valid to within a factor of 2 (in the worst case).
@@
There are indications of positive correlation between DM and scattering time (Fig. 5), and as the scattering time gets higher DM also increases and vice versa (the Pearson correlation coefficient 0.7 and the p-value ). However, we note that this correlation may be slightly exaggerated by the earlier-mentioned impact of scattering on timing-based DM measurements. A similarly strong correlation of scatter broadening time and DM (correlation coefficient 0.560.01) based on normal pulses was observed in Crab pulsar at 610 MHz by 2018MNRAS.479.4216M (61), and at 350 MHz by 2019MNRAS.483.1224D (22). Hence, our measurements are consistent with the earlier observations of scatter broadening and DM correlations at higher frequencies. Nevertheless, we note that more data and robust DM measurements methods are required to confirm the correlation in the presence of strong scatter broadening.
@@
The lag correlation of DM and scatter broadening time variations yielded a lag consistent with zero (to within a daily cadence of our observations). This is consistent with the results reported by 2018MNRAS.479.4216M (61), but significantly differs from the 30 days lag observed by 2008A&A…483…13K (45). It seems that the DM/ increase event observed by 2008A&A…483…13K (45) was caused by a plasma screen with a different structure and/or geometry of the Earth, pulsar, screen.
@@
Finally, we have also compared scatter broadening times with earlier observations. Thus, the measured scatter broadening times were scaled to 610 MHz with a scaling factor as measured for the Crab pulsar 2019ApJ…874..179K (44, 11, 63). The resulting distribution is very similar and only slightly shifted towards lower scatter-broadening times with respect to the distribution in Figure 3 in 2018MNRAS.479.4216M (61). This indicates that the scattering conditions were likely similar during both events, and possibly caused by similar plasma structures.
@@
One of the most interesting aspects originally noticed during the first observations in December 2024, were large variations in the numbers of detected GPs between 1-hour observations on different days (i.e. the rate of detectable GPs varied). During the first night of observations (2024-12-14) about 1400 GPs were detected by PRESTO, while on 2024-12-25 it reached 3000 to drop down to around 250 in the middle of February 2025 (Fig. 4). Interestingly, it started to increase again at the end of February 2025 to reach 2000 at the end of March (on 2025-03-27). The number of detected GPs was strongly correlated with the scatter broadening time (compare Figures 3 and 4, and see correlation in Figure 6).
The number of detected GPs as a function of scattering time has been modelled (black curve in Figure 6) using the scatter broadening time dependent fluence distribution (the fitted linear function in Figure 7), and functional form of pulse profile. Equation 15 in A was used to model giant pulse profiles as a function of time for different values of scatter broadening time with other parameters fixed (, , and ms). As discussed in Section 5.3 the fluence distribution of GPs depends on scatter broadening time. Hence, we used the linear dependence as fitted in Figure 7 to calculate normalisation of the fluence distribution for a specific . Then the expected number of GPs for particular value of was calculated as an integral of the fluence distribution above the minimal fluence . This minimal fluence was calculated as the fluence corresponding to the SNR threshold () used by the detection software PRESTO (SNR=5, i.e. peak flux density = 5230 Jy ms). The number of GPs as a function of and the results of our modelling are shown as a black curve in Figure 6. We note that this is only an approximation, because as discussed earlier by using convolution with boxcar function, PRESTO effectively implements a fluence detector, and this has not been reflected in our model. However, even this simple model reproduces the observed dependence very well. This gives us confidence that the variations in number of detected GPs can be explained by the up and down variations of the fluence distribution caused by refractive scintillation, which is also supported by the weeks timescale of these variations. Similar, variations in GP rates were observed by 1995ApJ…453..433L (54) in their largest ever recorded sample of GPs, which they also attributed to propagation effects. The explanation in terms of refractive scintillation is discussed in Section 6.1.2.
@@ 5.4 Flux density, SNR and spectral density distributions
Using the calibrated peak flux density and the distance to the pulsar D=2 kpc 2023ApJ…952..161L (51, 38), the peak spectral luminosity was calculated according to the following equation:
where is the peak flux density of the pulse. The resulting distribution is shown in the right panel of Figure 8. The maximum values are Lerg/s/Hz, which is consistent with the Crab spectral luminosities reported in literature (e.g. compare with Fig. 3 in 2022NatAs…6..393N (66) or Fig. 1 in 2018NatAs…2..865K (43)). These maximum spectral luminosities are a 2 – 3 orders of magnitude lower than the weakest repeating FRBs. However, assuming that there is enough energy in the system, the wait time for erg/s/Hz pulse is relatively short (of the order of 50,000 years). Hence, extra-galactic Crab-like pulsars may be responsible for weak repeating FRBs from the local Universe (see discussion in Section 6.2.1). However, the observed spectral luminosities are more than 10 orders of magnitude lower than typical or brightest FRBs. Hence, it is highly unlikely that such FRBs can be produced by objects similar to Crab pulsar (see discussion in Section 6.2) as there is not enough energy available. Additionally, based on our spectral luminosity distribution, the waiting time for such bright pulses (erg/s/Hz) is more than 10 orders of magnitude longer than the Hubble time.
@@ 6.1 Modelling DM and variations
We consider two distances between Earth and the screen (): (i) standard case midway (1000 pc) between the Earth and pulsar, and (ii) close to the pulsar (some 1900 pc from Earth) and consistent with the Crab Nebula or more broadly Perseus arm of the galaxy (explanation will be provided in Section 6.1.2). These two cases correspond to the ratio x= (Fig. 10) equal to 1 for (i) and 20 for (ii) respectively. For x=1 the velocity of the scintillation pattern with respect to observer on Earth is approximately equal to the velocity of the pulsar (eq. 10), while it can be higher for due to “lever-arm” argument (see equation 10 or discussion in 1995ApJ…451..717G (30)).
@@ 6.1.1 Modelling DM variations
The characteristic spatial scale of the plasma blob resulting from the fit is pc, and pc is consistent with the cloud size R pc estimated above using the mean values. This value leads to peak electron density 2600 e-cm-3. As an additional check, we performed a standard DM structure function analysis (for description see Section 5.2 in 2018MNRAS.479.4216M (61)) and obtained a characteristic timescale of 12.9 days, which corresponds to pc (when multiplied by 120 km/s) and agrees nearly perfectly with the value.
Hence, all tested methods lead to consistent results, and the resulting parameters are physically plausible. Despite the different epochs of observations, frequencies etc., they are also in-line with the studies by 2018MNRAS.479.4216M (61, 45, 29) as summarised in Table 1. This suggests that similar scintillation/scattering events due to plasma screens of the size pc and electron densities of the order of – are relatively common for the Crab pulsar. The observed electron densities are consistent with the filamentary structures inside the Crab Nebula (e.g. 1957PASP…69..227O (69, 26)).
@@
The baseline level of scattering time in this work was around 0.1 ms which corresponds to 2.5 ms at 215 MHz. However, sudden drops down to 0 ms were also observed (see Figure 2 in 2018MNRAS.479.4216M (61)). Since we are not sure if these were valid measurements of extremely low scattering time we used the baseline level as a reference.
@@ 6.1.2 Effects of refractive scintillation
The long observed timescale 70 days of the variations in GP rate and fluence distribution is a manifestation of intensity modulations caused by refractive scintillation. Moreover, estimates of the size pc of the plasma cloud agree very well with the large-scale structures required by refractive scintillation (see, for example 1999ApJ…514..272B (9)). The refractive scintillation effects on Crab GPs and average profiles have been observed and discussed before (e.g. 1995ApJ…453..433L (54), 1990MNRAS.244…68R (74), and 2011MNRAS.410..499G (29)). Interestingly, in a recent paper, 2024ApJ…973…87D (21) discuss the possibility of using GP rate as an ISM probe, which is exactly what transpired from our observations.
Under this hypothesis and using our equation 9 and equation 4.45 from 2012hpa..book…..L (53) (see also equation 4.1 in 1986MNRAS.220…19R (75)), we obtain:
@@
Additionally, the scatter broadening time is related to by the equation (e.g. 2012puas.book…..L (57)) :
@@
Typically, scintillation analysis (e.g. 2014ApJ…791L..32B (8)), uses auto-correlation function (ACF) of dynamic spectrum containing scintles to calculate scintillation time and de-correlation bandwidth as respectively the full-width at 1/e and half-maximum half-width of the 2D Gaussian fitted to the correlation peak in ACF. In this analysis the fluence normalisation or number of GPs were sampled daily in 1 hour observations. Hence, long-timescale dynamic spectrum is not available, and we can only use 1D temporal dependence of the number of GPs or normalisation of the fluence distribution (). Hence, in order to be consistent with this convention, we estimate by fitting Gaussian to the peak of the normalisation of the fluence distribution (not shown here but very similar to Figure 4) as a function of time, and calculate 51 days as full width at 1/e of the peak.
In order to calculate using equation 8, the distance to the screen and are also required. The additional constraints are imposed by equation 9 (above) and the following equation (equation 3 in 2014ApJ…791L..32B (8)):
@@
The first solution 0.05 pc could potentially be related to the inner edge of the solar system (distance 0.01 pc), but this does not have convincing physical explanation. On the other hand, the second solution (1835 pc) places the screen close to the pulsar, and the distance matches the distance to the Perseus spiral arm of the Milky Way galaxy, where the Crab pulsar is located. Hence, the scattering screen could be placed somewhere at the edge of the Perseus arm as seen along the LoS from Earth, which was also proposed by earlier studies (e.g. 1971ApJ…166..513C (18)). Alternatively, given the accuracy of this simplistic approach, and uncertainties of the distance to the pulsar, velocities of the screen and pulsar it is possible that the scattering material corresponds to the expanding Crab Nebula surrounding the pulsar. We note that solar elongations of Crab pulsar during the observing period were between 180°and 74°- sufficiently large to make any contributions from the solar wind negligible.